OpenStudy (anonymous):

Find the area enclosed by the polar curve r= 2-cos (theta)

OpenStudy (mrnood):

\[Area = \int\limits_{0}^{2\pi} (2-\cos \theta ) d \theta \]

OpenStudy (anonymous):

@MrNood it's \[Area=\int\limits_{0}^{2\pi} \frac{ 1 }{ 2 } f(\theta)^2 \] isn't it?

OpenStudy (anonymous):

forgot to tag \[d \theta \] at the end

OpenStudy (mrnood):

Not quite sure where that's coming from. I was working on integral r dtheta

OpenStudy (mrnood):

I'm a bit rusty on this stuff - so happy to be overridden!

OpenStudy (anonymous):

I got an answer of 4pi, so I think that's right

OpenStudy (mrnood):

seems unlikely value if you think of the function. In fact the answer MUST include r since the area obviously depends on r If the radius was constant then the area would be pi r^2 In this function the radius decreases at points around the polar sweep so I would expect the area to be < pi r^2 As I say - a little rusty but...

OpenStudy (mrnood):

the plot looks something like: